# Yongtao Li, Lihua Feng, Yuejian Peng

### v1

@@ 7 Concluding remarks
Finding the limit superior of $C$ such that
$t(G)\geq Cn^{2}\left(\lambda-\frac{n}{2}\right).$
Remark. By the clique density theorem of Reiher [49], we know that if $\varepsilon>0$ and $G$ is a graph on $n$ vertices with $e(G)\geq(\frac{1}{4}+\varepsilon)n^{2}$, then $G$ has at least $(\frac{1}{2}+o(1))\varepsilon n^{3}$ triangles, where $o(1)\to 0$ whenever $\varepsilon\to 0$. With aid of this powerful result, one can easily prove that if $G$ has at least $(\frac{1}{4}+\varepsilon)n^{2}$ edges, where we set $\varepsilon=\frac{1}{2}(\frac{\lambda}{n}-\frac{1}{2})$, then $G$ has at least $(\frac{1}{4}+o(1))n^{2}(\lambda-\frac{n}{2})$ triangles. Moreover, the graph $T_{n,2,1}$ shows that the coefficient $\frac{1}{4}$ is achievable. Very recently, Shengtong Zhang told us that taking $G=T_{n,3}$ yields $C\leq\frac{2}{9}$.
Another question as to how many triangles appear at least in any $n$-vertex graph with given spectral radius turns out to be interesting. For instance, for any graph $G$ with $\lambda(G)=\gamma n$, where $\gamma\in(\frac{1}{2},\frac{2}{3}]$, that is, $\lambda(T_{n,2})<\lambda(G)\leq\lambda(T_{n,3})$, there might exist a graph, say $L$, whose shape is approximately between $T_{n,2}$ and $T_{n,3}$, and $\lambda(L)=\gamma n$ such that $t(L)\leq t(G)$. Formally, we now define the minimizer $L$ as below. For $s\in\mathbb{N}$ and $\alpha>0$, we define $L_{n,s,\alpha}$ as a complete multipartite graph with $V(L_{n,s,\alpha})=V_{1}\cup\cdots V_{s}\cup V_{s+1}$, where $|V_{1}|=\cdots=|V_{s}|=\frac{n(1+\alpha)}{s+1}$ and $|V_{s+1}|=\frac{n(1-s\alpha)}{s+1}$. Next, we conjecture that $L_{n,s,\alpha}$ contains the minimum number of triangles among all graphs with given spectral radius.
@@
Nosal [47] proved that every graph $G$ with $m$ edges and $\lambda(G)>\sqrt{m}$ contains a triangle; see, e.g., [38, 44]. Recently, Ning and Zhai [45] showed that if $\lambda(G)\geq\sqrt{m}$, then $t(G)\geq\lfloor\frac{\sqrt{m}-1}{2}\rfloor$, unless $G$ is a complete bipartite graph. We propose the following conjecture.
@@
Recall in Theorem 1.2 that for every $q<n/2$, the graph $G$ with $\lfloor n^{2}/4\rfloor+q$ edges contains at least $q\lfloor n/2\rfloor$ triangles. While, in our Theorem 2.2, we proved the spectral version for $n\geq 300q^{2}$. Under our framework, it seems possible to improve this bound slightly. For the sake of clarity, we do not make any serious attempt to optimize the absolute constant in our statements and proofs. It is ambitious to extend our result to the following wide range.
@@
Whether Theorems 2.2 and 2.3 hold for every $q<n/2$? If not, what is the limit superior of $\delta$ such that the spectral results hold for $q<\delta n$?
In 2017, Pikhurko and Yilma [48, Theorem 1.4] proved that for every color-critical graph $F$ with $\chi(F)=r+1$, there exist $\delta=\delta(F)>0$ and $n_{0}$ such that for all $n\geq n_{0}$ and $q<\delta n$, if $G$ is an $n$-vertex graph with $e(T_{n,r})+q$ edges and it has the minimum number of copies of $F$, then $G$ is obtained from $T_{n,r}$ by adding $q$ new edges. Theorem 1.2 gives $\delta(K_{3})=1/2$. More generally, a result of Lovász and Simonovits [34, 35] implies that $\delta(K_{r+1})=1/r$ for every $r\geq 2$. Moreover, Pikhurko and Yilma [48] also determined the limit superior of $\delta$ for a number of color-critical graphs, such as, odd cycles $\delta(C_{2k+1})=1/2$, and cliques with one edge removed $\delta(K_{r+2}-e)=(r-1)/r^{2}$. In a forthcoming paper, we will use a variety of different methods to establish a spectral counting result for any color-critical graph. As the techniques are quite complicated, we intend to present it separately $\heartsuit$.
Declaration of competing interest
The authors declare that they have no conflicts of interest to this work.
Data availability
No data was used for the research described in the article.
@@ Acknowledgements
The authors are thankful to Xiaocong He and Loujun Yu for carefully reading an early manuscript of this paper. Yongtao Li would like to thank Dheer Noal Desai for the inspiring discussions on spectral graph problems over the years. Li was supported by the Postdoctoral Fellowship Program of CPSF (No. GZC20233196), Feng was supported by the NSFC grant (No. 12271527), and Peng was supported by the NSFC grant (No. 11931002).

### v2

@@ 7 Concluding remarks
What is the best constant $C>0$ such that $t(G)\geq Cn^{2}\left(\lambda-\frac{n}{2}\right)$?
Remark. By the clique density theorem of Reiher [52], we know that if $\varepsilon>0$ and $G$ is a graph on $n$ vertices with $e(G)\geq(\frac{1}{4}+\varepsilon)n^{2}$, then $G$ has at least $(\frac{1}{2}+o(1))\varepsilon n^{3}$ triangles, where $o(1)\to 0$ whenever $\varepsilon\to 0$. With the aid of this powerful result, one can easily prove that if $G$ has at least $(\frac{1}{4}+\varepsilon)n^{2}$ edges, where we set $\varepsilon:=\frac{1}{2}(\frac{\lambda}{n}-\frac{1}{2})$, then $G$ has at least $(\frac{1}{4}+o(1))n^{2}(\lambda-\frac{n}{2})$ triangles. Moreover, the graph $T_{n,2,1}$ shows that the coefficient $\frac{1}{4}$ is achievable. Recently, Shengtong Zhang told us that taking $G=T_{n,3}$ yields $C\leq\frac{2}{9}$.
Another question as to how many triangles appear at least in any $n$-vertex graph with given spectral radius turns out to be interesting. For instance, for any graph $G$ with $\lambda(G)=\gamma n$, where $\gamma\in(\frac{1}{2},\frac{2}{3}]$, that is, $\lambda(T_{n,2})<\lambda(G)\leq\lambda(T_{n,3})$, there might exist a graph, say $L$, whose shape is approximately between $T_{n,2}$ and $T_{n,3}$ such that $\lambda(L)=\gamma n$ and $t(L)\leq t(G)$. Formally, we now define the minimizer $L$ as follows. For $s\in\mathbb{N}$ and $\alpha>0$, we define $L_{n,s,\alpha}$ as a complete multipartite graph with $V(L_{n,s,\alpha})=V_{1}\cup\cdots V_{s}\cup V_{s+1}$, where $|V_{1}|=\cdots=|V_{s}|=\frac{n(1+\alpha)}{s+1}$ and $|V_{s+1}|=\frac{n(1-s\alpha)}{s+1}$. Next, we speculate that $L_{n,s,\alpha}$ contains the minimum number of triangles among all graphs with given spectral radius.
@@
Nosal [50] proved that every graph $G$ with $m$ edges and $\lambda(G)>\sqrt{m}$ contains a triangle; see, e.g., [41, 47]. In 2023, Ning and Zhai [48] showed that if $\lambda(G)\geq\sqrt{m}$, then $t(G)\geq\lfloor\frac{\sqrt{m}-1}{2}\rfloor$, unless $G$ is a complete bipartite graph. We propose the following conjecture.
@@
Theorem 1.2 asserts that every graph $G$ with $\lfloor n^{2}/4\rfloor+q$ edges contains at least $q\lfloor n/2\rfloor$ triangles for every $q<n/2$. In Theorem 2.2, we proved a spectral version for all $q\leq\frac{1}{11}\sqrt{n}$. This bound on $q$ is tight up to a constant factor (Example 3.3). Under our framework, it seems feasible to improve the constant factor slightly. However, new ideas are needed to determine the exact bound on $q$. It is ambitious to propose the following problem.
@@
What is the supremum $\delta\in(0,0.8)$ such that Theorem 2.2 holds for every integer $q<\delta\sqrt{n}$? Additionally, what is the optimal range of $q$ in Theorem 2.3?
In 2017, Pikhurko and Yilma [51] proved that for every color-critical graph $F$ with $\chi(F)=r+1$, there exist $\delta=\delta(F)>0$ and $n_{0}$ such that for all $n\geq n_{0}$ and $q<\delta n$, if $G$ is an $n$-vertex graph with $e(T_{n,r})+q$ edges and it achieves the minimum number of copies of $F$, then $G$ is obtained from the Turán graph $T_{n,r}$ by adding $q$ new edges. Furthermore, by Theorem 1.2, we see that $\delta(K_{3})=\frac{1}{2}$. A result of Lovász and Simonovits [37, 38] implies that $\delta(K_{r+1})=\frac{1}{r}$ for every $r\geq 2$. Moreover, Pikhurko and Yilma [51] also determined the exact value of $\delta$ for a number of color-critical graphs, such as, odd cycles $\delta(C_{2k+1})=\frac{1}{2}$, and cliques with one edge removed $\delta(K_{r+2}-e)=\frac{r-1}{r^{2}}$. Similarly, it is also meaningful to determine the exact value of $\delta(F)$ for various graphs $F$ in the spectral setting.
@@ Acknowledgements
We are grateful to Xiaocong He and Loujun Yu for carefully reading an early draft of the paper. Yongtao Li would like to thank Dheer Noal Desai for inspiring discussions on spectral graph problems. Lihua Feng was supported by the National Natural Science Foundation of China (Nos. 12271527 and 12471022), and Yuejian Peng was supported by the National Natural Science Foundation of Hunan Province (No. 2025JJ30003) and the National Natural Science Foundation of China (No. 12571363).